# ZTF-EP-24-09

### v1

@@
Abstract. During the era of NASA’s Apollo missions Keith S. Runcorn proposed an explanation of discrepancy between the Moon’s negligible global magnetic field and magnetised samples of lunar regolith, based on identical vanishing of external magnetic field of a spherical shell, magnetized by an internal source which is no longer present. We revisit and generalize the Runcorn’s result, showing that it is a consequence of a (weighted) orthogonality of gradients of harmonic functions on a spherical shell in arbitrary number of dimensions. Furthermore, we explore bounds on external magnetic field in the case when the idealized spherical shell is replaced with a more realistic geometric shape and when the thermoremanent magnetization susceptibility deviates from the spherical symmetry. Finally, we analyse a model of thermoremanent magnetization acquired by crustal inward cooling of a spherical astrophysical body and put some general bounds on the associated magnetic field.
@@ 1 Introduction
Magnetic field permeates all matter in Solar System and offers us an important insight into the history and internal structure of the astrophysical bodies. Above all, here we encounter solar and planetary dynamos, movement of electrically conducting fluid described by magnetohydrodynamics. Also, even if a terrestrial planet or a moon no longer has an active dynamo in its core (or, indeed, never had one), it still may possess an imprint of the past magnetic fields in a form of the remanent magnetization in its outer, solid part.
@@
A particularly intriguing and instructive episode revolves around Earth’s natural satellite, Moon. Early investigations of lunar magnetism were performed with orbital probes Luna 2 [DYZP61] in 1959, Luna 10 [DYZP66] in 1966 and Explorer 35 [SCC67] in 1967, none of which detected global lunar magnetic field above the sensitivity of their magnetometers, with the order of magnitude ranging between 10 nT and 0.1 nT. These observations were later corroborated by orbital measurements within missions Apollo 15 in 1971 and Apollo 16 in 1972 (cf. [Ful74, Rus80, HVF91] and references therein), putting an upper bound on lunar magnetic dipole moment 8 orders of magnitude below the present Earth’s magnetic dipole moment. On the other hand, analysis of the returned samples in the Apollo missions have revealed that lunar rocks possess natural remanent magnetization, carried by iron particles, providing us with abundant evidence for the past lunar magnetic field.
@@
Keith S. Runcorn [Run75a, Run75b, Run77] has proposed an explanation of the seemingly inconsistent lunar magnetism [^2] An overview of theories for the origin of lunar magnetism proposed during 1970s is given in [DD79]., based on the following simple, lucid observation. If a spherical shell bears a permanent magnetization proportional to a magnetic field, sourced by currents contained within the inner sphere, than the resulting magnetic field out of the outer sphere, in the absence of the source currents, is exactly zero. Relying on data obtained from several seismometers, placed on lunar surface during Apollo missions [Wie09] we know that Moon is a differentiated body, with crust, mantle and core. Runcorn’s scenario suggests that Moon once had active dynamo in its core, whose magnetic field has magnetized lunar crust; as the lunar dynamo is no longer active, external magnetic field of the remanent crustal magnetization is zero. Unfortunately, as was discussed already by Runcorn himself [Run75b, Run77], this model offers only a partial solution of the puzzle. First of all, Moon is not ideally spherical, albeit one does not expect dramatic change of the conclusion as the Moon’s deviation from spherical symmetry is of the order of $10^{-3}$ [Run77]. Secondly, although Runcorn allows radially dependent proportionality between magnetization and core’s magnetic field, it does not take into account that outer layers will additionally magnetize inner layers (consequence of which was investigated by Srnka [Srn76a, Srn76b]). Overall, these amendments will results with some limited global external magnetic field. However, further intricacies of the problem proliferate with localised magnetic fields, usually referred to as “magnetic anomalies”, reaching the order of hundreds of nT. These were originally observed in Apollo missions, with orbital and in situ measurements, and more recently mapped in detail by Lunar Prospector mission [MHL+08] (for a discussion about the possible relation between lunar magnetic anomalies and impact processes cf. [WWS12, OWS+20, W+23, YW24]). Finally, although the consensus in the community seems to be that Moon once had long-lived dynamo [WT14, OWS+20, TE22, W+23] this conclusion was put under question with more recent reassessment of the lunar samples [T+21]. Nevertheless, it is essential to stress that even if the Runcorn’s basic argument is insufficient to explain lunar magnetic field, it still remains a valuable idea which might be applied to some other astrophysical bodies, as was previously shown in the case of the Hermean [Ste76a] and Martian [LS97] magnetic fields.
@@
Looking back at the original Runcorn’s theorem we are lead to a number of glaring questions. Where does this result “come from”, namely can it be reformulated as a corollary to some more general property of an appropriately chosen family of functions? What happens with the conclusions once the geometry of the problem is “slightly perturbed”? What can be deduced from a model which takes into account magnetization between the layers of the crust? We shall examine these questions more closely and provide some insights which could be useful for future investigations of the crustal magnetism. The paper is organized as follows. In section 2 we state and prove a generalized Runcorn’s theorem, a property of harmonic functions. In section 3 we analyse magnetic field of a magnetized crust in the cases when some of the Runcorn’s assumptions are not met. In section 4 we revisit Srnka’s model of a crust, magnetized during the inward cooling, and look more closely into the pertaining bounds on the magnetic field.
@@ 2 Generalized Runcorn’s theorem
More concretely, let $\Omega$ be a spherical shell (“crust” of the astrophysical body), bounded by concentric spheres of radii $0<a<b$, while matter was magnetized with field $\bm{H}_{0}$, sourced by currents contained in ball $r<a$ (which are no longer present). As the exterior $r>b$ is simply connected and, by assumption, devoid of free currents, we may write $\bm{H}_{0}=-\nabla\Psi_{0}$, with scalar potential $\Psi_{0}$ which is a solution of the Laplace’s equation, $\Delta\Psi_{0}=0$, and vanishes at infinity. Crucial physical assumption is that magnetization is linear in the sense that $\bm{M}=\kappa\bm{H}_{0}$ with some function $\kappa$, that is
@@
Let $\mathscr{U}=\{\bm{x}\in\mathds{R}^{m}\mid\left\lVert\bm{x}\right\rVert<r_{+}\}$ and $\mathscr{V}=\{\bm{x}\in\mathds{R}^{m}\mid\left\lVert\bm{x}\right\rVert>r_{-}\}$ be open sets, with $m\geq 2$ and radii $0<r_{-}<r_{+}$. Furthermore, let $u$ be a harmonic function on $\mathscr{U}$, $v$ a harmonic function on $\mathscr{V}$, vanishing at infinity, $\lim_{\left\lVert\bm{x}\right\rVert\to\infty}v(\bm{x})=0$, and $f:\left<0,\infty\right>\to\mathds{R}$ a $C^{1}$ function. Then for any $r_{-}<a<b<r_{+}$ and a spherical shell $\mathcal{A}=\{\bm{x}\in\mathds{R}^{m}\mid a\leq\left\lVert\bm{x}\right\rVert\leq b\}$ the following orthogonality holds,
@@
where $p_{k}$ and $q_{\ell}$ are harmonic polynomials of degree, respectfully, $k$ and $\ell$. These series converge absolutely and uniformly on any compact subset of $\mathcal{A}$, allowing us to perform term by term derivations and the exchange of integrals with the sums.
In preparation for the evaluation of the integrals we stress several elementary observations. Each $\bm{x}\neq\bm{0}$ can be decomposed as $\bm{x}=r\bm{s}$, with $\bm{s}\in S\mathrel{\mathop{:}}=\{\bm{y}\in\mathds{R}^{m}\mid\left\lVert\bm{y}\right\rVert=1\}$. Homogeneity of the harmonic polynomials implies that $p_{k}(\bm{x})=p_{k}(r\bm{s})=r^{k}p_{k}(\bm{s})$ and $\partial_{r}p_{k}(\bm{x})=\partial_{r}(r^{k}p_{k}(\bm{s}))=kr^{k-1}p_{k}(\bm{s})$. Furthermore, if we denote by $\mathrm{d}\sigma$ surface element on the unit sphere $S$, than the surface and the volume elements may be decomposed, respectfully, as $\mathrm{d}a=r^{m-1}\,\mathrm{d}\sigma$ and $\mathrm{d}V=r^{m-1}\mathrm{d}r\,\mathrm{d}\sigma$. Finally, harmonic polynomials are orthogonal on the unit sphere $S$ in the sense that (see e.g. proposition 5.9 in [ABR01])
@@
One might ask whether a converse of the Theorem 2.1 holds: If a function $f$ satisfies condition (7) for all harmonic functions $u$ and $v$, satisfying conditions from the theorem, is $f$ necessarily isotropic, $f=f(\left\lVert\bm{x}\right\rVert)$? Discussion in the Appendix B of [AHD96] suggests that the answer is positive but, as far as we are aware of, there is no rigorous proof of this conjecture in the literature. Closely related question arises in the context of the inverse problems [LV20], reconstruction of the magnetization from the measured external field (cf. [PN10] for an analysis of the Lunar Prospector data). This problem is more general in a sense that $\Omega$ does not have to be a spherical shell and magnetization does not have to be a multiple of gradient of harmonic function, but it is more special in a sense that one of the harmonic functions is fundamental solution for the Laplacian. One of the major obstacles here is the nonuniqueness which stems from the existence of “magnetic annihilators” [MH03], also referred to as “invisible magnetizations” [BGKM21], domains $\Omega$ with magnetizations $\bm{M}$ with magnetic field which vanishes in the whole or some of the connected components of the complement of $\Omega$.
@@ 3 Approximate Runcorn’s theorem
Now we turn to various deviations from the initial setting of the Runcorn’s theorem, when some of the geometric or analytic assumptions are relaxed. A direct step towards generalization could start with the pointwise Cauchy–Bunyakowsky–Schwartz inequality,
@@
for all $p,q,r\in\left[1,\infty\right]$ such that $p^{-1}+q^{-1}+r^{-1}=1$, assuming that all terms on the right hand side are well-defined. However, from a physical point of view, it is more useful to look directly at the magnetic field, as an upper bound on the value of the scalar potential is not telling us much about its gradient. For simplicity (and clarity) we shall focus on the $m=3$ case only.
In the complement of the magnetized domain $\Omega$, for all $\bm{r}\in\mathds{R}^{3}-\overline{\Omega}$, field $\bm{H}$ may be written directly from the equation (5), as
@@
with $\bm{u}\mathrel{\mathop{:}}=(\bm{r}-\bm{r}^{\prime})/\left\lVert\bm{r}-\bm{r}^{\prime}\right\rVert$. There are two basic problems, that we may first treat separately: the one in which the function $\kappa$ is no longer isotropic and the one in which the crust is “slightly deformed” spherical shell.
@@ 3.1 Anisotropic susceptibility
where we have introduced the abbreviation $V(\mathcal{A})=4\pi(b^{3}-a^{3})/3$ for the volume of the spherical shell $\mathcal{A}$. Now, for any $\alpha>2$ we have
@@
with the equality holding at the given point iff $\bm{H}_{0}$ and $\bm{u}$ are linearly dependent at that point. In our case we have $\alpha=3$, so that
@@
All together, for any $r>b/\delta$ we have
@@
Given that one can estimate the volume $V(\mathcal{A})$, e.g. from seismographic data, and magnitude $H_{0}$ of the magnetizing field, e.g. from the rock samples, inequality (21) offers, at least in principle, a bound between the measured magnetic field $\bm{B}$ and the measure of the susceptibility’s anisotropy $\varepsilon$ (in the case of the Moon, some rough estimates of the orders of the magnitude may be inferred from the seismographic data [Wie09] and measurements from the lunar rock samples [WT14]). As long as $\delta<1/2$, corresponding to distances at least as twice the radius of the body, we have $(1-\delta)^{-3}<8$, so that this factor increases constant $C$, no more than one order of magnitude. However, as we approach the surface of the astrophysical body, the upper bound becomes less useful as the limit $\delta\to 1$ implies $C\to\infty$. An alternative approach is to use the formula, derived from (4),
@@ 3.2 Nonspherical crust
Let us now assume that the susceptibility $\kappa$ is isotropic, $\kappa=\kappa(r)$, while the crust is “slightly deformed” in a sense that the domain $\Omega$ is bounded by two smooth disjoint embedded surfaces $\mathcal{S}_{-}$ and $\mathcal{S}_{+}$, each homeomorphic to the 2-sphere, and admits embedding of a spherical shell $\mathcal{A}_{0}=\{\bm{x}\in\mathds{R}^{3}\mid a\leq\left\lVert\bm{x}\right\rVert\leq b\}\subseteq\Omega$, as sketched in the Figure 1. For the sake of better estimate, it is advantageous to choose spherical shell of volume as large as possible. We shall denote by $\Omega^{\prime}\mathrel{\mathop{:}}=\Omega-\mathcal{A}_{0}$ part of the domain which remains after we remove the spherical shell $\mathcal{A}_{0}$.
$\mathcal{A}_{0}$
@@
[…]

### v2

@@
Abstract. During the era of NASA’s Apollo missions, Keith S. Runcorn proposed an explanation of discrepancy between the Moon’s negligible global magnetic field and magnetized samples of lunar regolith, based on identical vanishing of external magnetic field of a spherical shell, magnetized by an internal source which is no longer present. We revisit and generalize the Runcorn’s result, showing that it is a consequence of a (weighted) orthogonality of gradients of harmonic functions on a spherical shell in arbitrary number of dimensions. Furthermore, we explore bounds on external magnetic field in the case when the idealized spherical shell is replaced with a more realistic geometric shape and when the thermoremanent magnetization susceptibility deviates from the spherical symmetry. Finally, we analyse a model of thermoremanent magnetization acquired by crustal inward cooling of a spherical astrophysical body and put some general bounds on the associated magnetic field.
@@ 1 Introduction
Magnetic field permeates all matter in Solar System, offering important insight into the history and internal structure of astrophysical bodies. Above all, here we encounter solar and planetary dynamos, movement of electrically conducting fluid described by magnetohydrodynamics. Also, even if a terrestrial planet or a moon no longer has an active dynamo in its core (or, indeed, never had one), it still may possess an imprint of the past magnetic fields in a form of the remanent magnetization in its outer, solid layers.
@@
A particularly intriguing and instructive episode revolves around Earth’s natural satellite, the Moon. Early investigations of lunar magnetism were conducted with orbital probes Luna 2 [DYZP61] in 1959, Luna 10 [DYZP66] in 1966 and Explorer 35 [SCC67] in 1967, none of which detected global lunar magnetic field above the sensitivity of their magnetometers, with the order of magnitude ranging between 10 nT and 0.1 nT. These observations were later corroborated by orbital measurements from the Apollo 15 mission in 1971 and the Apollo 16 mission in 1972 (cf. [Ful74, Rus80, HVF91] and references therein), putting an upper bound on lunar magnetic dipole moment 8 orders of magnitude below the present Earth’s magnetic dipole moment. On the other hand, analysis of the samples returned by the Apollo missions have revealed that lunar rocks possess natural remanent magnetization, carried by iron particles, providing us with abundant evidence for the past lunar magnetic field.
@@
Keith S. Runcorn [Run75a, Run75b, Run77] has proposed an explanation of the seemingly inconsistent lunar magnetism [^2] An overview of theories for the origin of lunar magnetism proposed during 1970s is given in [DD79]., based on the following simple, lucid observation. If a spherical shell bears a permanent magnetization proportional to a magnetic field, sourced by currents contained within the inner sphere, then the resulting magnetic field out of the outer sphere, in the absence of the source currents, is exactly zero. Based on data from several seismometers, placed on lunar surface during Apollo missions [Wie09], we know that the Moon is a differentiated body, with crust, mantle and core. Runcorn’s scenario suggests that the Moon once had active dynamo in its core, whose magnetic field has magnetized lunar crust; as the lunar dynamo is no longer active, external magnetic field of the remanent crustal magnetization is zero. Unfortunately, as was discussed already by Runcorn himself [Run75b, Run77], this model offers only a partial solution of the puzzle. First of all, the Moon is not ideally spherical, albeit one does not expect dramatic change of the conclusion as the Moon’s deviation from spherical symmetry is of the order of $10^{-3}$ [Run77]. Secondly, although Runcorn allows for a radially dependent proportionality between magnetization and core’s magnetic field, it does not take into account that outer layers will additionally magnetize inner layers (consequence of which was investigated by Srnka [Srn76a, Srn76b]). Overall, these amendments result in a limited global external magnetic field. However, further intricacies of the problem proliferate with localised magnetic fields, usually referred to as “magnetic anomalies”, reaching the order of hundreds of nT. These were originally observed in Apollo missions, through orbital and in situ measurements, and more recently mapped in detail by Lunar Prospector mission [MHL+08] (for a discussion about the possible relation between lunar magnetic anomalies and impact processes cf. [WWS12, OWS+20, W+23, YW24]). Finally, although the consensus in the community seems to be that the Moon once had long-lived dynamo [WT14, OWS+20, TE22, W+23], this conclusion was put under question with more recent reassessment of the lunar samples [T+21]. Nevertheless, it is essential to stress that even if the Runcorn’s basic argument is insufficient to explain lunar magnetic field, it still remains a valuable idea which might be applied to some other astrophysical bodies, as was previously shown in the case of the Hermean [^1] Modern consensus, based on all available data (cf. [MHM+18] and references therein), is that a core dynamo is favoured over remanent crustal magnetization as the main candidate source of the Mercury’s magnetic field. [Ste76a] and Martian [LS97] magnetic fields.
@@
Looking back at the original Runcorn’s theorem, we are led to a number of glaring questions. Where does this result “come from”, namely can it be reformulated as a corollary to some more general property of an appropriately chosen family of functions? What happens with the conclusions once the geometry of the problem is “slightly perturbed”? What can be deduced from a model which takes into account magnetization between the layers of the crust? We shall examine these questions more closely and provide some insights which could be useful for future investigations of the crustal magnetism. The paper is organized as follows. In section 2 we state and prove a generalized Runcorn’s theorem, a property of harmonic functions. In section 3 we analyse magnetic field of a magnetized crust in the cases when some of the Runcorn’s assumptions are not met. In section 4 we revisit Srnka’s model of a crust, magnetized during the inward cooling, and look more closely into the pertaining bounds on the magnetic field.
@@ 2 Generalized Runcorn’s theorem
More concretely, let $\Omega$ be a spherical shell (“crust” of the astrophysical body), bounded by concentric spheres of radii $0<a<b$, while matter was magnetized with field $\bm{H}_{0}$, sourced by currents contained in ball $r<a$ (which are no longer present). As the exterior $r>b$ is simply connected and, by assumption, devoid of free currents, we may write $\bm{H}_{0}=-\nabla\Psi_{0}$, with scalar potential $\Psi_{0}$ which is a solution of the Laplace’s equation, $\Delta\Psi_{0}=0$, and vanishes at infinity. Crucial physical assumption is that magnetization is linear in the sense that $\bm{M}=\kappa\bm{H}_{0}$ with some function $\kappa$, that is,
@@
Let $\mathscr{U}=\{\bm{x}\in\mathds{R}^{m}\mid\left\lVert\bm{x}\right\rVert<r_{+}\}$ and $\mathscr{V}=\{\bm{x}\in\mathds{R}^{m}\mid\left\lVert\bm{x}\right\rVert>r_{-}\}$ be open sets, with $m\geq 2$ and radii $0<r_{-}<r_{+}$. Furthermore, let $u$ be a harmonic function on $\mathscr{U}$, $v$ a harmonic function on $\mathscr{V}$, vanishing at infinity, $\lim_{\left\lVert\bm{x}\right\rVert\to\infty}v(\bm{x})=0$, and $f:\left<0,\infty\right>\to\mathds{R}$ a $C^{1}$ function. Then for any $r_{-}<a<b<r_{+}$ and a spherical shell $\mathcal{A}=\{\bm{x}\in\mathds{R}^{m}\mid a\leq\left\lVert\bm{x}\right\rVert\leq b\}$, the following orthogonality holds:
@@
where $p_{k}$ and $q_{\ell}$ are harmonic polynomials of degree, respectfully, $k$ and $\ell$. These series converge absolutely and uniformly on any compact subset of $\mathcal{A}$, allowing us to perform term-by-term derivations and the exchange of integrals with the sums.
In preparation for the evaluation of the integrals, we stress several elementary observations. Each $\bm{x}\neq\bm{0}$ can be decomposed as $\bm{x}=r\bm{s}$, with $\bm{s}\in S\mathrel{\mathop{:}}=\{\bm{y}\in\mathds{R}^{m}\mid\left\lVert\bm{y}\right\rVert=1\}$. Homogeneity of the harmonic polynomials implies that $p_{k}(\bm{x})=p_{k}(r\bm{s})=r^{k}p_{k}(\bm{s})$ and $\partial_{r}p_{k}(\bm{x})=\partial_{r}(r^{k}p_{k}(\bm{s}))=kr^{k-1}p_{k}(\bm{s})$. Furthermore, if we denote by $\mathrm{d}\sigma$ surface element on the unit sphere $S$, then the surface and the volume elements may be decomposed, respectfully, as $\mathrm{d}a=r^{m-1}\,\mathrm{d}\sigma$ and $\mathrm{d}V=r^{m-1}\mathrm{d}r\,\mathrm{d}\sigma$. Finally, harmonic polynomials are orthogonal on the unit sphere $S$ in the sense that (see e.g. proposition 5.9 in [ABR01])
@@
One might ask whether a converse of the Theorem 2.1 holds: If a function $f$ satisfies condition (7) for all harmonic functions $u$ and $v$, satisfying conditions from the theorem, is $f$ necessarily isotropic, $f=f(\left\lVert\bm{x}\right\rVert)$? Discussion in Appendix B of [AHD96] suggests that the answer is positive but, as far as we are aware of, there is no rigorous proof of this conjecture in the literature. Closely related question arises in the context of the inverse problems [LV20], reconstruction of the magnetization from the measured external field (cf. [PN10] for an analysis of the Lunar Prospector data). This problem is more general in a sense that $\Omega$ does not have to be a spherical shell and magnetization does not have to be a multiple of gradient of harmonic function, but it is more special in a sense that one of the harmonic functions is fundamental solution for the Laplacian. One of the major obstacles here is the nonuniqueness which stems from the existence of “magnetic annihilators” [MH03], also referred to as “invisible magnetizations” [BGKM21], domains $\Omega$ with magnetizations $\bm{M}$ with magnetic field which vanishes in the whole or some of the connected components of the complement of $\Omega$.
@@ 3 Approximate Runcorn’s theorem
Now, we turn to various deviations from the initial setting of the Runcorn’s theorem, when some of the geometric or analytic assumptions are relaxed. A direct step towards generalization could start with the pointwise Cauchy–Bunyakovsky–Schwarz inequality,
@@
for all $p,q,r\in\left[1,\infty\right]$ such that $p^{-1}+q^{-1}+r^{-1}=1$, assuming that all terms on the right-hand side are well-defined. However, from a physical point of view, it is more useful to look directly at the magnetic field, as an upper bound on the value of the scalar potential is not telling us much about its gradient. For simplicity (and clarity), we shall focus on the $m=3$ case only.
In the complement of the magnetized domain $\Omega$, for all $\bm{r}\in\mathds{R}^{3}-\overline{\Omega}$, field $\bm{H}$ may be written directly from equation (5), as
@@
with $\bm{u}\mathrel{\mathop{:}}=(\bm{r}-\bm{r}^{\prime})/\left\lVert\bm{r}-\bm{r}^{\prime}\right\rVert$. There are two basic problems that we may first treat separately: the one in which the function $\kappa$ is no longer isotropic and the one in which the crust is “slightly deformed” spherical shell.
@@ 3.1 Anisotropic susceptibility
where we have introduced the abbreviation $V(\mathcal{A})=4\pi(b^{3}-a^{3})/3$ for the volume of the spherical shell $\mathcal{A}$. Now, for any $\alpha>2$, we have
@@
with the equality holding at the given point iff $\bm{H}_{0}$ and $\bm{u}$ are linearly dependent at that point. In our case, we have $\alpha=3$, so that
@@
Altogether, for any $r>b/\delta$, we have
@@
As long as $\delta<1/2$, corresponding to distances at least as twice the radius of the body, we have $(1-\delta)^{-3}<8$, so that this factor increases constant $C$ no more than one order of magnitude. However, as we approach the surface of the astrophysical body, the upper bound becomes less useful as the limit $\delta\to 1$ implies $C\to\infty$. An alternative approach is to use the formula, derived from (4),
@@ 3.2 Nonspherical crust
Let us now assume that the susceptibility $\kappa$ is isotropic, $\kappa=\kappa(r)$, while the crust is “slightly deformed” in a sense that the domain $\Omega$ is bounded by two smooth disjoint embedded surfaces $\mathcal{S}_{-}$ and $\mathcal{S}_{+}$, each homeomorphic to the 2-sphere, and admits embedding of a spherical shell $\mathcal{A}_{0}=\{\bm{x}\in\mathds{R}^{3}\mid a\leq\left\lVert\bm{x}\right\rVert\leq b\}\subseteq\Omega$, as sketched in Figure 1. For the sake of better estimate, it is advantageous to choose spherical shell of volume as large as possible. We shall denote by $\Omega^{\prime}\mathrel{\mathop{:}}=\Omega-\mathcal{A}_{0}$ part of the domain which remains after we remove the spherical shell $\mathcal{A}_{0}$.
@@
where $V(\Omega^{\prime})$ is the volume of the domain $\Omega^{\prime}$. An in-depth analysis of magnetic field for the particular case of ellipsoidal bodies was performed in [JWL99].